# Learning a Stable, Safe, Distributed Feedback Controller for a Heterogeneous Platoon of Vehicles

### v1

@@ Abstract
Platooning of autonomous vehicles has the potential to increase safety and fuel efficiency on highways. The goal of platooning is to have each vehicle drive at some speed (set by the leader) while maintaining a safe distance from its neighbors. Many prior works have analyzed various controllers for platooning, most commonly linear feedback and distributed model predictive controllers. In this work, we introduce an algorithm for learning a stable, safe, distributed controller for a heterogeneous platoon. Our algorithm relies on recent developments in learning neural network stability and safety certificates. We train a controller for autonomous platooning in simulation and evaluate its performance on hardware with a platoon of four F1Tenth vehicles. We then perform further analysis in simulation with a platoon of 100 vehicles. Experimental results demonstrate the practicality of the algorithm and the learned controller by comparing the performance of the neural network controller to linear feedback and distributed model predictive controllers.
I INTRODUCTION
@@
In safety-critical systems where performance guarantees are of paramount importance, adoption of learning-based controllers has understandably been slow. This is because most learning-based controllers, up until recently, have not had the ability to provide performance guarantees over the range of scenarios the system can expect to encounter. Recent work in learning controllers, safety certificates, and stability certificates has opened the door for learning safe and reliable controllers for safety-critical systems [1]. As we consider using these controllers in real systems, however, developing algorithms that will learn controllers that translate well from the simulated model to the real world is critical. Further, for these methods to work on multi-agent systems, it will be necessary to develop algorithms that scale well as the number of agents increases.
@@
Autonomous vehicle platooning is an automated driving technology that can improve safety [2] and reduce fuel emissions [3] on our roads. However, there is an inherent tradeoff between these two objectives as fuel efficiency is maximized when vehicles are closer together while safety is maximized when vehicles are further apart. This tradeoff between performance (a stable, tightly packed platoon) and safety (the vehicles do not crash) makes the platooning problem a good case study for certifiable learning-based control.
@@
In this work, we focus on learning decentralized feedback controllers for a heterogeneous platoon of vehicles. The focus is on developing controllers that are provably stable for the entire platoon (learn a centralized stability certificate based on a decentralized controller). The difficulty in learning a stable decentralized controller for a multi-agent system where we need centralized performance guarantees arises from the fact that the computational complexity of verifying stability grows exponentially as the size of the system grows. This is because verifying stability requires solving mixed-integer linear programs (MILPs).
The contributions of this work are as follows: 1) we introduce an algorithm for learning a certifiably stable neural network controller that enables the engineer to “guide” the resulting controller to desired behavior; 2) we use a change of variable to reformulate the dynamics of a heterogeneous platoon as that of a homogeneous platoon, enabling us to learn and verify a single controller for a heterogeneous platoon; 3) we train controllers in simulation and validate the learned controllers on hardware using four F1Tenth vehicles [4], shown in Fig. 1, and in simulation with 100 vehicles.
II RELATED WORK
II-A Autonomous vehicle platooning
@@
In [7], the authors showed that using a linear feedback controller and a CDH spacing policy can lead to string instability, meaning disturbances witnessed earlier in the platoon will propagate and become worse toward the end of the platoon. Further, they showed how the disturbance to error gain increases as the size of the platoon increases for both a predecessor following (PF) and bidirectional (BD) topology. Building on this, the authors of [8] showed how the stability margins of a platoon using a linear feedback controller with BD topology decays to zero as $O(1/N^{2})$ when each vehicle has the same dynamics. This work extended [9], where the authors also showed how a particular nonlinear controller outperformed linear controllers for a platoon using a PF topology.
@@
If vehicles can communicate with each other, distributed model predictive control (DMPC), which requires vehicles share their planned trajectories with their neighbors, is possible. One of the first DMPC controllers was analyzed in [13] where it was shown that a DMPC controller is both stable and string stable for a PF topology when all of the vehicles in the platoon know the desired velocity of the platoon a priori. In [14], the stability condition was extended for the case that not all vehicles know the desired velocity a priori. In [15], two DMPC controllers, both slight deviations from the controller proposed in [14], were validated on both hardware and in simulation. The simulation results, which analyzed the performance of DMPC as the size of the platoon scaled up to 100 vehicles, showed that DMPC greatly outperforms linear feedback controllers.
II-B Certifiable learning-based control
@@
Upon analyzing much of the prior theoretical results in platoon control, we arrive at the question: can we learn better controllers for platooning? Specifically, since many vehicles on the road today are equipped with forward-facing radar and adaptive cruise control technology, it would be beneficial to learn controllers for CDH spacing policies and PF topologies that close the performance gap between linear feedback and DMPC, thus enabling tightly-packed platoons and greater traffic capacity. To achieve this goal, we need to ensure these learned controllers are safe and stable, which is where the emerging research areas of learning-based control [16] and neural network verification [17, 18] merge into certifiable learning-based control. To date, research within certifiable learning-based control has generally been related to either reachability analysis [19] or learning Lyapunov and barrier certificates [1].
@@
In this work, we are interested in learning Lyapunov functions to guarantee stability of a platoon using a distributed neural network controller. One of the earlier works in certifiable learning-based control was [20], where both a neural Lyapunov function and a controller were learned to prove stability when controlling a given nonlinear system. In [21], this was extended to consider also systems with learned neural network dynamics. Depending on the system being analyzed, it can be straightforward to extend these earlier works directly to multi-agent systems, but the certification problem quickly becomes intractable as the number of agents (and thus the state size) increases due to the need to solve a nonconvex mixed integer program. In [22], the authors get around this due to the fact that the goal of each agent is independent of others, and thus stability certificates can be learned in a completely decentralized manner.
III METHODOLOGY
III-A Controller design
@@
We consider a platoon of $N$ vehicles indexed by $i=1,\ldots,N$ where vehicle 1 is the leader. Each vehicle has dynamics given by
$s_{i}(k+1)=\begin{bmatrix}1&\Delta t\,\\ 0&1-\frac{\Delta t\,}{\tau_{i}}\end{bmatrix}s_{i}(k)+\begin{bmatrix}0\\ \frac{\Delta t\,}{\tau_{i}}\end{bmatrix}a_{i}(k)$ (1)
@@
where $s_{i}(k)=(p_{i}(k),v_{i}(k))\in\mathop{\mathbb{R}}^{2}$ and $a_{i}(k)\in\mathop{\mathbb{R}}$ are vehicle $i$’s state (position and velocity) and control input (desired velocity), respectively, $\tau_{i}$ is a vehicle longitudinal dynamics delay parameter, and $\Delta t\,$ is the discrete timestep. It is common in the literature to use a three-state linear dynamics model with position, velocity, and acceleration as the state and desired acceleration as the control input, similar to [23, 8, 14]. However, we use Eq. 1 for two reasons: 1) it models the F1Tenth vehicle’s longitudinal dynamics well and 2) it uses desired velocity instead of desired acceleration as the control input, similar to the F1Tenth vehicles.
@@
We assume each vehicle in the platoon can sense the relative position and velocity of the vehicle in front of it. Our goal is to learn a distributed control law $\pi:\mathop{\mathbb{R}}^{2}\to\mathop{\mathbb{R}}$ such that each vehicle $i$ drives at the same speed as its predecessor $i-1$ while maintaining some predefined distance $d_{i,i-1}$. The controller is distributed in the sense that each vehicle runs its controller locally based on information it receives from its local sensors. The input to the controller is the error state
@@
$u_{i}(k)=\frac{a_{i}(k)-v_{i}(k)}{\tau_{i}},\quad i=1,\ldots,N.$ (3)
With this change of variable, the dynamics for vehicle $i$ becomes discrete-time double integrator dynamics given by
@@
$A=\begin{bmatrix}1&\Delta t\,\\ 0&1\end{bmatrix}\quad B=\begin{bmatrix}0\\ \Delta t\,\end{bmatrix}.$
We note that without this change of variable, it would be considerably more difficult to verify a Lyapunov function for the platoon over the range of possible $\tau_{i}$ parameters due to the coupling between the state variables and the dynamics parameters. Based on Eq. 4 and the definition of the error state $x_{i}$, it is not hard to show that the error dynamics of the entire platoon is given by
@@
These dynamics are equivalent to a string of double integrators [9], thus enabling us to use various insights from the relevant literature to guide our control design. For instance, in [9] it was shown that a nonlinear PF controller that saturates performs better than a linear controller. Thus, we can be confident that placing bounds on the control input (which would always be required in practice) can lead to better performance.
III-B Controller learning and verification
Our goal is to learn a neural network controller $\pi:\mathop{\mathbb{R}}^{2}\to\mathop{\mathbb{R}}$ and a Lyapunov function $V:\mathop{\mathbb{R}}^{2N}\to\mathop{\mathbb{R}}$ such that the platoon is exponentially stable within some compact region ${\cal X}$ containing the origin. From Lyapunov theory, this is equivalent to
@@ III-B Controller learning and verification
Equation 6a is trivially satisfied by setting the bias terms of the feedforward neural network $V$ to zero. For neural network verification, we cannot have strict inequalities in the conditions we would like to verify. Thus we convert the condition Eq. 6b to the condition $V(x)\geq\varepsilon_{1}\|x\|_{1}$. Now suppose we have learned a controller and a Lyapunov function and we would like to verify the conditions Eqs. 6b and 6c are true. As described in [21], this can be achieved by solving the optimization problems
@@
$\max_{x\in{\cal X}}{\cal L}_{\text{dec}}(x)=V(\bar{A}x+\bar{B}\pi(x))-(1+\varepsilon_{2})V(x)$ (7b)
and verifying their solutions are equal to zero. Note that the two objective functions measure the violations of the Lyapunov conditions given by Eqs. 6b and 6c. These two optimization problems can be encoded as mixed-integer linear programs (MILPs) and solved to global optimality. We do this using CVXPY [24] and Gurobi [25].
@@
Data: Initial $V$, $\pi$, dynamics $f$, ${\cal D}=\emptyset$, X, U, simulation environment env, small $\varepsilon$
@@ 7             $x\leftarrow f(x,\pi(x))$
Once the inner loop converges (when the system state converges to the origin), the Lyapunov conditions are checked. If satisfied, we have a learned controller that is certifiably stable for the system. If not, the process repeats. Optionally, we can train on the dataset ${\cal D}$ using stochastic gradient descent (or any other optimizer) to further update $V$ and $\pi$. We have found that augmenting the dataset with random points within ${\cal X}$ can help speed up convergence. Another option, as described in [21], is to augment ${\cal D}$ with any points that violate the Lyapunov conditions during the branching process of the MILP solver. Overall, algorithm 1 is very flexible and easily adjusted to incorporate ideas from other works.
IV EXPERIMENTS
IV-A Controller learning
Algorithm 1 returns a feedback controller $\pi:\mathop{\mathbb{R}}^{2}\to\mathop{\mathbb{R}}$. For each vehicle $i$, the controller $\pi$ takes in the vehicle’s current error state $x_{i}$, given by Eq. 2, and outputs the vehicle’s desired acceleration $u_{i}$, i.e., $u_{i}=\pi(x_{i})$. We are able to adjust the characteristics of this controller based on how we design ${\cal L}_{\pi}$. However, we were only able to verify a learned controller for up to $N=3$. Figure Fig. 2 shows the values of the solutions to Eqs. 7a and 7b for $N=2$ and $N=3$ while running algorithm 1.
For the experiment shown in Fig. 2, we parameterized $V$ and $\pi$ using feedforward neural networks with two hidden layers, eight neurons in each hidden layer, and leaky ReLU activation functions with negative slope of 0.1. In general, we have found leaky ReLU neural networks perform better with this algorithm than ReLU neural networks. We see that $N=2$ converges in 25 episodes whereas $N=3$ converges in 150 episodes. Each time both conditions are equal to zero, we increase the size of the error bounds ${\cal X}$ until we reach the error bounds we are interested in verifying. This explains why for both $N=2$ and $N=3$, both conditions reach zero multiple times, but each time except the last are actually for some $\tilde{\cal X}\subset{\cal X}$.
Figure 2: Per episode results for the maximum Lyapunov conditions given by Eq. 7a and Eq. 7b when running algorithm 1. Results are shown only for $N=2$ (convergence at episode 30) and $N=3$ (convergence at episode 150); convergence occurred after 8 episodes for $N=1$.
After $N=3$, we were never able to achieve convergence even when using a very small ${\cal X}=[-0.1,0.1]^{Nn}$. We allowed the algorithm to time out after 1000 episodes, which took roughly 18 hours on a computer with an Intel Core i7 and an Nvidia GeForce RTX 3080. Most of this time is spent trying to solve the MILP given by Eq. 7b or training on the dataset ${\cal D}$ for 300 epochs. Despite the fact we are unable to verify the controllers for larger $N$, we have found that the learned controller does still perform well, and it does not take very many episodes for the controller to begin performing well (in only a few episodes for small $N$ and after a few tens of episodes for larger $N$). Thus, for larger $N$, we ignore the verification step and run our algorithm to directly learn a controller.

### v2

# https://robot.neu.edu/

@@ Abstract
Platooning of autonomous vehicles has the potential to increase safety and fuel efficiency on highways. The goal of platooning is to have each vehicle drive at a specified speed (set by the leader) while maintaining a safe distance from its neighbors. Many prior works have analyzed various controllers for platooning, most commonly linear feedback and distributed model predictive controllers. In this work, we introduce an algorithm for learning a stable, safe, distributed controller for a heterogeneous platoon. Our algorithm relies on recent developments in learning neural network stability certificates. We train a controller for autonomous platooning in simulation and evaluate its performance on hardware with a platoon of four F1Tenth vehicles. We then perform further analysis in simulation with a platoon of 100 vehicles. Experimental results demonstrate the practicality of the algorithm and the learned controller by comparing the performance of the neural network controller to linear feedback and distributed model predictive controllers.
1 Introduction
@@
In safety-critical systems where performance guarantees are of paramount importance, adoption of learning-based controllers has understandably been slow. This is because most learning-based controllers, up until recently, have not had the ability to provide performance guarantees over the range of scenarios the system can expect to encounter. Recent work in learning controllers, safety certificates, and stability certificates has opened the door for learning reliable controllers for safety-critical systems [4]. As we consider using these controllers in real systems, however, developing algorithms that will learn controllers that translate well from simulation to the real world is critical. Furthermore, for these methods to work on multi-agent systems, it will be necessary to develop algorithms that scale as the number of agents increases.
@@
Autonomous vehicle platooning is an emerging driving technology that can improve safety [17] and reduce fuel emissions [14] on our roads. However, there is an inherent tradeoff between these two objectives as fuel efficiency is maximized when vehicles are closer together while safety is maximized when vehicles are further apart. This tradeoff between performance and safety makes the platooning problem a good case study for certifiable learning-based control.
In this work, we focus on learning decentralized feedback controllers for a heterogeneous (i.e., each vehicle can have different dynamics) platoon of vehicles. The focus is on developing controllers that are provably stable for the entire platoon (learn a centralized stability certificate based on a decentralized controller). The difficulty in learning a stable decentralized controller for a multi-agent system where we need centralized performance guarantees arises from the fact that the computational complexity of verifying stability grows exponentially as the size of the system (number of vehicles) grows. This is because verifying stability requires solving mixed-integer linear programs (MILPs).
The contributions of this work are as follows: 1) we introduce an algorithm for learning a certifiably stable neural network controller that enables the engineer to “guide” the resulting controller to desired behavior; 2) we use a change of variable to reformulate the dynamics of a heterogeneous platoon as that of a homogeneous platoon, enabling us to learn and verify a single controller for a heterogeneous platoon; 3) we train controllers in simulation and validate the learned controllers on real hardware using four F1Tenth vehicles [19], shown in Fig. 1, and in simulation with 100 vehicles.
@@
2 Related Work
2.1 Autonomous vehicle platooning
@@
In [21], the authors showed that using a linear feedback controller and a CDH spacing policy can lead to string instability, meaning disturbances witnessed earlier in the platoon will propagate and become worse toward the end of the platoon. Additionally, they showed how the disturbance to error gain increases as the size of the platoon increases for both a predecessor following (PF) and bidirectional (BD) topology. Building on this, the authors of [28] showed how the stability margins of a platoon using a linear feedback controller with BD topology decays to zero as $O(1/N^{2})$ when each vehicle has the same dynamics. This work extended [12], where the authors also showed how a particular nonlinear controller outperformed linear controllers for a platoon using a PF topology.
@@
If vehicles can communicate with one another, distributed model predictive control (DMPC), which requires vehicles share their planned trajectories with their neighbors, is possible. One of the first DMPC controllers was analyzed in [7] where it was shown that a DMPC controller is both stable and string stable for a PF topology with a CDH policy when all of the vehicles in the platoon know the desired velocity of the platoon a priori. In [29], the stability condition was extended for the case that not all vehicles know the desired velocity a priori. Additionally, the work of [22] extended that of [29] for the case of CTH topologies and arbitrary communication topologies. In [23], two DMPC controllers, both slight deviations from the controller proposed in [29], were validated on both hardware and in simulation. The simulation results, which analyzed the performance of DMPC as the size of the platoon scaled up to 100 vehicles, showed that DMPC greatly outperforms linear feedback controllers.
2.2 Certifiable learning-based control
@@
Upon analyzing much of the prior theoretical results in platoon control, we arrive at the question: can we learn better controllers for platooning? Specifically, since many vehicles on the road today are equipped with forward-facing radar and adaptive cruise control technology, it would be beneficial to learn controllers for CDH spacing policies and PF topologies that close the performance gap between linear feedback and DMPC, thus enabling tightly-packed platoons and greater traffic capacity. To achieve this goal, we need to ensure these learned controllers are stable, which is where the emerging research areas of learning-based control [1] and neural network verification [15, 8] merge into certifiable learning-based control. To date, research within certifiable learning-based control has generally been related to either reachability analysis [9] or learning Lyapunov and barrier certificates [4].
@@
In this work, we are interested in learning the parameters of a neural network Lyapunov function to guarantee stability of a platoon using a distributed neural network controller. One of the earlier works in certifiable learning-based control was [2], where both a neural Lyapunov function and a controller were learned to prove stability when controlling a given nonlinear system. In [3], this was extended to consider also systems with learned neural network dynamics. Depending on the system being analyzed, it can be straightforward to extend these earlier works directly to multi-agent systems, but the certification problem quickly becomes intractable as the number of agents (and thus the state size) increases due to the need to solve a nonconvex mixed integer program. In [20], the authors get around this due to the fact that the goal of each agent is independent of other agents, and thus stability certificates can be learned in a completely decentralized manner.
3 Methodology
3.1 Controller design
@@
We consider a platoon of $N$ vehicles indexed by $i=1,\ldots,N$ where vehicle 1 is the leader and $i<j$ implies vehicle $i$ precedes vehicle $j$ in the platoon. Each vehicle has dynamics given by
$s_{i}(k+1)=\begin{bmatrix}1&\Delta t\,\\ 0&1-\frac{\Delta t\,}{\tau_{i}}\end{bmatrix}s_{i}(k)+\begin{bmatrix}0\\ \frac{\Delta t\,}{\tau_{i}}\end{bmatrix}v^{\text{des}}_{i}(k)$ (1)
@@
where $s_{i}(k)=(p_{i}(k),v_{i}(k))\in\mathop{\mathbb{R}}^{2}$ and $v^{\text{des}}_{i}(k)\in\mathop{\mathbb{R}}$ are vehicle $i$’s state (position and velocity) and control input (desired velocity), respectively, $\tau_{i}$ is a vehicle longitudinal dynamics delay parameter, and $\Delta t\,$ is the discrete timestep. It is common in the literature to use a three-state linear dynamics model with position, velocity, and acceleration as the state and desired acceleration as the control input, similar to [24, 28, 29]. However, we use Eq. 1 for two reasons: 1) it models the F1Tenth vehicle’s longitudinal dynamics well and 2) it uses desired velocity instead of desired acceleration as the control input, similar to the F1Tenth vehicles.
@@
We assume each vehicle in the platoon can sense the relative position and velocity of the vehicle in front of it. Our goal is to learn a distributed control law $\pi:\mathop{\mathbb{R}}^{2}\to\mathop{\mathbb{R}}$ such that each vehicle $i$ drives at the same speed as its predecessor while maintaining some predefined distance $d_{i,i-1}$. The controller is distributed in the sense that each vehicle runs its controller locally based on information it receives from its sensors. The input to the controller is the error state
@@
$u_{i}(k)=\frac{v^{\text{des}}_{i}(k)-v_{i}(k)}{\tau_{i}},\quad i=1,\ldots,N.$ (3)
With this change of variable, the dynamics for vehicle $i$ become discrete-time double integrator dynamics given by
@@
$A=\begin{bmatrix}1&\Delta t\,\\ 0&1\end{bmatrix}\quad B=\begin{bmatrix}0\\ \Delta t\,\end{bmatrix}$
where the control input $u_{i}\in\mathop{\mathbb{R}}$ is the acceleration. We note that without this change of variable, it would be considerably more difficult to verify a Lyapunov function for the platoon over the range of possible $\tau_{i}$ parameters due to the coupling between the state variables and the dynamics parameters. Based on Eq. 4 and the definition of the error state $x_{i}$, it is not hard to show that the error dynamics of the entire platoon is given by
@@
Given Eq. 5, it is tempting to use optimal control techniques, e.g., a linear quadratic regulator (LQR). However, these techniques rely on each vehicle having complete knowledge of the control actions $u$ for the entire platoon, which is not practical for a distributed, real-time implementation.
Alternatively, the dynamics given by Eq. 5 are equivalent to a string of double integrators [12], thus enabling the use of various insights from the relevant literature to guide control design. For instance, in [12] it was shown that a nonlinear PF controller that saturates performs better than a linear controller. Thus, we can be confident that placing bounds on the control input (which would always be required in practice) can lead to better performance. Furthermore, this motivates the use of learned neural network controllers which can be more expressive than classical nonlinear controllers.
3.2 Controller learning and verification
Our goal is to learn a neural network controller $\pi:\mathop{\mathbb{R}}^{2}\to\mathop{\mathbb{R}}$ and Lyapunov function $V:\mathop{\mathbb{R}}^{2N}\to\mathop{\mathbb{R}}$ such that the platoon is exponentially stable within some compact region ${\cal X}$ containing the origin. Note that we can bound the control inputs by appending extra layers to the output layer of the neural network (exact formulation is provided in Section 6.1). From Lyapunov theory, exponential stability within a compact region ${\cal X}$ is guaranteed if
@@ 3.2 Controller learning and verification
Condition (6a) is trivially satisfied by setting the bias terms of the neural network $V$ to zero. For neural network verification, we cannot have strict inequalities in the conditions we would like to verify. Thus we convert the condition (6b) to the condition $V(x)\geq\varepsilon_{1}\|x\|_{1}$. Now suppose we have learned a controller $\pi$ and a Lyapunov function $V$ and we would like to verify the conditions (6b) and (6c) are true. This can be achieved by solving the optimization problems
@@
$\max_{x\in{\cal X}}{\cal L}_{\text{dec}}(x)=V(\bar{A}x+\bar{B}\pi(x))-(1-\varepsilon_{2})V(x)$ (7b)
and verifying their solutions are equal to zero. Note that the objective functions ${\cal L}_{\text{pos}}$ and ${\cal L}_{\text{dec}}$ measure the violations of the positivity Lyapunov condition (6b) and decreasing Lyapunov condition (6c), respectively. These two optimization problems can be encoded as mixed-integer linear programs (MILPs) by using mixed-integer formulations of the $\ell_{1}$-norm and the activation functions within the Lyapunov function and control policy (see Sections 6.2 and 6.3). Thus, the optimization problems given by Eqs. 7a and 7b can be solved to global optimality. We do this using CVXPY [6] and Gurobi [11].
@@
Once the inner loop converges, the Lyapunov conditions are checked. If satisfied, we have a learned controller that is certifiably stable for the system. If not, the process repeats. Optionally, we can train on the dataset ${\cal D}$ using stochastic gradient descent (or any other optimizer) to further update $V$ and $\pi$. We have found that augmenting the dataset with random points within ${\cal X}$ can help speed up convergence. Another option, as described in [3], is to augment ${\cal D}$ with any points that violate the Lyapunov conditions during the branching process of the MILP solver. Overall, algorithm 1 is very flexible and easily adjusted to incorporate ideas from other works.
Data: Initial $V$, $\pi$, dynamics $f$, ${\cal D}=\emptyset$, ${\cal X}$, simulation environment env, $\varepsilon$
@@ 7             $x\leftarrow f(x,\pi(x))$
4 Experiments
4.1 Controller learning
Algorithm 1 returns a feedback controller $\pi:\mathop{\mathbb{R}}^{2}\to\mathop{\mathbb{R}}$. For each vehicle $i$, the controller $\pi$ takes in the vehicle’s current error state $x_{i}$, given by Eq. 2, and outputs the vehicle’s desired acceleration $u_{i}$, i.e., $u_{i}=\pi(x_{i})$. We are able to adjust the characteristics of this controller based on how we design ${\cal L}_{\pi}$. However, we were only able to verify a learned controller for up to $N=3$. Figure 2 shows the values of the solutions to Eqs. 7a and 7b for $N=2$ and $N=3$ while running algorithm 1.
For the experiment shown in Fig. 2, we parameterized $V$ and $\pi$ using feedforward neural networks with two hidden layers, eight neurons in each hidden layer, and leaky ReLU activation functions with a slope of 0.1. In general, we have found that when using algorithm 1, neural networks with leaky ReLU activation functions converge more easily than neural networks with ReLU activation functions. We see that with $N=2$, we achieve convergence in about 30 episodes whereas with $N=3$, we achieve convergence in about 150 episodes. Each time both conditions are equal to zero, we increase the size of the error bounds ${\cal X}$ until we reach the error bounds we are interested in verifying. This explains why for both $N=2$ and $N=3$, both conditions reach zero multiple times, but each time except the last are actually for some $\tilde{\cal X}\subset{\cal X}$.
Figure 2: Per episode results for the maximum Lyapunov condition violations given by Eq. 7a and Eq. 7b when running algorithm 1. Results are shown only for $N=2$ (convergence at episode 30) and $N=3$ (convergence at episode 150); convergence occurred after 8 episodes for $N=1$.
After $N=3$, we were unable able to achieve convergence even when using a very small ${\cal X}=[-0.1,0.1]^{Nn}$. We allowed the algorithm to time out after 1000 episodes, which took roughly 18 hours on a computer with an Intel Core i7 and an Nvidia GeForce RTX 3080. Most of this time is spent trying to solve the MILP given by Eq. 7b or training on the dataset ${\cal D}$ for 300 epochs. Despite the fact that we are unable to verify the controllers for larger $N$, we have found that the learned controller does still perform well, and it does not take very many episodes for the controller to begin performing well (in only a few episodes for small $N$ and after a few tens of episodes for larger $N$). Thus, for larger $N$, we ignore the verification step and use algorithm 1 to directly learn a controller.